OpenStudy (anonymous):

What are the values of the functions y =x3 - 4 for x = 0, 1, 2 and 3? -4, -1, 2, 5 0 ,3 , 6, 9 -4, 2, 5, 8 0, 1, 5, 9

OpenStudy (anonymous):

just substitute and see waht you get

OpenStudy (whpalmer4):

Plug in the numbers and find out! However, it appears that you didn't copy the function correctly...

OpenStudy (anonymous):

substitute each of those values and tell me your answer

OpenStudy (anonymous):

I fixed it.

OpenStudy (whpalmer4):

sort of. If you mean \[y=x^3-4\]you should write y=x^3-4

OpenStudy (whpalmer4):

Anyhow, same procedure applies. \[y = (0)^3-4 = \]\[y = (1)^3-4=\]\[y=(2)^3-4 = \]\[y=(3)^3-4=\]

OpenStudy (whpalmer4):

No, it looks like you mean \[y = 3x-4\]

OpenStudy (anonymous):

-4 -3 4 23 and yes

OpenStudy (whpalmer4):

people copying and pasting material into openstudy often inadvertently convert expressions like \[y = x^3\] into y=x3, so most of us assume that an exponent was involved there. However, strictly speaking \[y=x3 \]is equivalent to \[y = 3x\]but one doesn't normally write the numeral after the variable. So, if we are agreed that \[y = 3x-4\] is the one true equation here :-) \[y = 3(0)-4 = \]\[y = 3(1)-4=\]\[y=3(2)-4 = \]\[y = 3(3)-4 = \]And you should find that set of numbers as one of your answer choices.